OpenStudy (anonymous):

g(x) = x2 +6x + 1 Find the solutions of g(x). Show each step. Can someone please help?

OpenStudy (campbell_st):

this needs the general quadratic equation \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a= 1, b = 6 and c = 1 substitute the values then simplify

OpenStudy (anonymous):

okay, one sec

OpenStudy (anonymous):

x=-6(sqrt)32/2

OpenStudy (anonymous):

where do i go from here?

OpenStudy (campbell_st):

well \[\sqrt{32} = 4\sqrt{2}\] so make the substitution and then divide by 2

OpenStudy (anonymous):

I'm really confused on this

OpenStudy (campbell_st):

ok... so using some laws for radicals \[\sqrt{32} = \sqrt{16 \times 2} = \sqrt{16} \times \sqrt{2} = 4\sqrt{2}\] so the problem is \[x = \frac{-6 \pm 4\sqrt{2}}{2}\] so divide -6 by 2 and then 4 by 2 and you get \[x = -3 \pm 2\sqrt{2}\]

OpenStudy (anonymous):

okay, thank you. i wasn't sure if i divided 4 by 2 as well

OpenStudy (anonymous):

where do i from here?

OpenStudy (campbell_st):

thats the answer the 2 values of x are \[x = - 3 + \sqrt{2}...and.... x = -3 - \sqrt{2}\]

OpenStudy (anonymous):

oh! okay thank you so much!

OpenStudy (campbell_st):

oops it I missed a bit \[x = -3 + 2\sqrt{2} .... and ..... x = -3 - 2\sqrt{2}\] sorry about that